# Quantum Computing for Automotive Applications

### v2

@@ 1 Introduction
However, recently the expectations have become more realistic. A 2024 BCG report [11] continued to predict significant economic impact but highlights challenges, notably the limitations of Noisy Intermediate-Scale Quantum (NISQ) hardware [12] and the slow progress in quantum algorithms. Meanwhile, advances in classical computing – especially in hardware like GPUs and AI algorithms – have significantly raised the performance bar.
@@ 3 Optimization
Another challenge is the mapping between the optimization problem and the physical quantum system. Often, no direct mapping exists and many mapping incur significant overheads [43]. Consequently, research focuses on developing these mappings, either focusing on specific hardware (e. g., neutral atom hardware [44]), or particular problems (knapsack problem [17], robot motion planning [45, 23]; see Table 2).
@@
Several quantum-inspired optimization strategies have emerged as byproducts of the research in quantum optimization. The marriage of generative modeling and quantum-inspired tensor network methods resulted in the generator-enhanced optimization (GEO) framework [49]. In GEO, a quantum-inspired generative model is used to propose new candidate solutions by identifying patterns in high-quality solutions. The performance of GEO has since been analyzed on the production planning use case, where it has been shown to perform on par with other conventionally used solvers [50].
@@ Discussion
It is still unclear whether quantum computing can provide significant value for real-world optimization workloads [52]. With large, noisy quantum devices, the key challenge is identifying hardware-native problem instances, e. g., maximum independent set problems for neutral atom hardware, that can be embedded with minimal overhead while remaining difficult for classical solvers [43, 53, 54]. While optimization of such problems on near-term quantum hardware devices has produced some impressive results [38], it is not yet feasible at the scales required for industrially relevant problems. However, with advancements in hardware and algorithms, scaling analyses suggest that quantum computing may eventually outperform classical methods for specific instances particularly suited to quantum devices [55, 56].
@@ 4 Simulation
HHL: The HHL algorithm uses quantum phase estimation (QPE) [83] to implicitly write the eigenvalues of matrix $A$ in a quantum register. It then stores the inverted eigenvalues in the quantum amplitudes, effectively inverting $A$. After reversing QPE and a postselection step, the algorithm prepares a state proportional to |$\ket{x}$, given that the input state was |$\ket{b}$. The overall complexity is $\mathcal{O}(\log(N)\kappa^{2}/\epsilon)$ [80], where $N$ corresponds to the dimension of $A$, $\kappa$ is its condition number quantifying the difficulty of inverting $A$, and $\epsilon$ is the precision of the solution. This is an exponential advantage in $N$ compared to the best-known classical methods. However, the promised speedup can be significantly reduced or even destroyed by considering the algorithmic overhead for state preparation, implementation of matrix oracles, or the readout of the full solution [84]. For example, in the case of FEM, the anticipated exponential speedup of the HHL algorithm is reduced to a polynomial speedup by considering the precision of the readout [85].
@@ 5 Materials Science and Quantum Chemistry
VQE: This type of algorithm starts with a variational circuit (VQC) ansatz, which attempts to prepare the system’s ground state. In an iterative process, the variational parameters of the VQC are adapted via gradient descent, trying to minimize the system’s energy [109, 107]. VQE is considered a near-term algorithm that is suitable for a NISQ computers. However, it suffers from trainability issues when large systems are considered or as the depth and complexity of the VQCs grow, see the barren plateau problem [77]. When a suitable set of parameters for the VQC is found, and a trial state is prepared, the expected value of the Hamiltonian can be estimated, and thus, its energy and eigenstate can be found.
@@
Qubitization or Block encoding: All quantum circuits are unitary transformations. In this sense, unitary operators are native to quantum computers. There is no native way to encode arbitrary operators, e. g., general Hamiltonians or arbitrary operators, in a quantum register and operate over them. Such ability is necessary for quantum simulations. To simulate Hamiltonian dynamics using a quantum computer it is necessary to use the technique of qubitization or block encoding [86], which is an FT algorithm primitive. The technique consists of encoding an arbitrary operator inside a larger unitary operator through auxiliary qubits. This technique is probabilistic, which adds to the sample complexity of the algorithms. When this is achieved, additional algorithms can perform computations with the encoded data, as we comment below.
@@ 6 Machine Learning
Progress in AI is largely driven by the scaling laws that describe the improvement of AI as a function of the progress in computing, data, and model parameters [120]. Similarly, Sutton’s ”bitter lesson” [121] emphasizes the long-term success of algorithms that leverage increased computational power over those relying on complex, domain-specific optimizations. As classical computing infrastructures start to hit scaling limits [122], quantum computing may eventually provide the computational capabilities needed to align with this principle.
Quantum machine learning (QML) [123, 124] integrates quantum algorithms with machine learning, aiming to reduce training times and enhancing the scalability and quality of models. While some QML algorithms promise theoretical improvements in computational complexity [125, 126, 127, 128], the practical advantages of quantum machine learning have yet to be conclusively established on large-scale industry-relevant datasets. In the following sections, we describe near-term, fault-tolerant, and quantum-inspired QML, concluding with a discussion on the limitations of QML.
@@
Representing classical data on a quantum computer: QML algorithms can be applied to both quantum and classical data. For classical data, input data must be mapped to a quantum state, using a procedure referred to as data or feature encoding. Notably, encoding methods that leverage entanglement, such as amplitude encoding, are particularly interesting because they allow for exponential data compression by encoding a feature vector of size $N$ into $\log_{2}N$ qubits. However, methods like amplitude encoding incur an exponential runtime in the number of qubits, and any potential quantum speed-up is lost at the encoding stage [129]. A promising solution lies in using approximations to address the challenge of exponential runtime in amplitude encoding. Jobst et al. [96] show that classical data with an algebraically decaying Fourier spectrum can be well-approximated by a quantum circuit with a linear number of nearest-neighbor two-qubit gates.
@@
Quantum kernel methods: Quantum kernel methods [130, 131] are one approach in QML for supervised learning task. They leverage quantum computing by replacing the classical kernel of a support vector machine (SVM) with a quantum kernel. In this approach, a classical data vector $\mathbf{x}$ is mapped into a high-dimensional feature space using a feature map that is represented by a parameterized quantum circuit $U(\mathbf{x})$. Quantum computers can efficiently evaluate the inner products between data points in this quantum feature space. By determining all pair-wise inner products of the embedded data points, the kernel matrix is generated. This kernel matrix is then used within the SVM framework to perform classification or regression tasks.
Supervised training of variational circuits: Supervised training of variational quantum circuits [132, 133, 134] (VQC) involves optimizing the parameters of a VQC to minimize a predefined cost function, typically derived from labeled data. The circuit processes input data $x$ encoded into quantum states by a feature map $U(x)$. This state is then evolved via a VQC $V(\theta)$, before the measurement is performed. The expectation value of the observable acts as the prediction of the model. The parameters $\theta$ are tuned during the training to minimize a cost function. A method for computing gradients on a quantum computer is the parameter-shift rule [132, 135].
Generative training of variational circuits: Quantum states prepared by VQCs naturally represent probability distributions that can be used to generate synthetic data. Therefore, VQCs can also be used for generative learning. Instead of fitting the output of the model to the data label, the output is fitted to a distribution. Two widely used training routines in quantum generative modeling, as depicted in Figure 3, are the quantum circuit Born machine (QCBM) [136, 137] and the quantum generative adversarial network (QGAN) [138, 139], as depicted in Figure 4. For a comprehensive review of these methods, see the referenced literature [140]. The QCBM aims to minimize a statistical distance such as the Kullback-Leibler (KL) divergence, between the probability distributions of the generated and target data by adjusting model parameters with gradient-based or gradient-free optimization techniques.
@@ 7 Benchmarking
Aggregate benchmarks evaluate a larger part of the quantum computing stack considering a broader range of attributes. An example is Quantum Volume (QV), which takes into account the number of qubits, the quality of gate operations, and the accuracy of measurements [173]. While the QV is a well-established benchmark, it has known drawbacks like its strong dependency on high-quality qubits, which favors technologies like ion-trapped quantum computers (which reach QV values of over a million [174, 175]) compared to superconducting hardware (currently reaching QV values well below 1000 [176]). Hence, other aggregate benchmarks like circuit layer operations per second (CLOPS) [177] or generalized volumetric benchmarks [178] based on QV address these shortcomings and offer alternatives.
@@ 8 Conclusion and Future Directions
Quantum Chemistry Fuel Cell [103] Quantinuum H1 using 8 qubits Simulation of Oxygen Reduction Reaction (ORR) on platinum/cobalt (Pt/Co) surface with active space of 4 electrons in 4 orbitals.
@@
Quantum hardware has significantly progressed. However, this progress has not matched initial industry expectations, particularly in scale. This limitation is evident from the data presented in Table 2, which summarizes our recent case studies. Our case studies utilized various quantum hardware platforms, e. g., approximately 5,000 qubits on D-Wave’s Advantage 4.1, 137 qubits on QuEra’s Aquilla, 11 qubits on IBM’s Kolkata, 8-10 qubits on Quantinuum’s H1 and IonQ’s Harmony machines.
@@
Future Directions: Hardware continues to advance, as evidenced by enhanced qubit fidelities [214], error correction breakthroughs [215, 216], and new distributed, multi-QPU architectures [217]. Emerging quantum-classical architectures, also referred to as HPCQC [218] or quantum-centric supercomputing [8], integrate Quantum Processing Units (QPUs) as accelerators into the high-performance computing stack. While these hardware trends are promising, they require adaptation to algorithms [219, 220]. These adaptations often involve techniques such as circuit cutting [221], hardware-software co-design [209], and error mitigation to optimize performance and reliability.

### v3

@@ 1 Introduction
However, recently, the expectations have become more realistic. A 2024 BCG report [11] continued to predict significant economic impact but highlights challenges, notably the limitations of Noisy Intermediate-Scale Quantum (NISQ) hardware [12] and the slow progress in quantum algorithms. Meanwhile, advances in classical computing – especially in hardware like GPUs and AI algorithms – have significantly raised the performance bar.
@@ 3 Optimization
Another challenge is mapping the optimization problem and the physical quantum system. Often, no direct mapping exists, and many mappings incur significant overheads [43]. Consequently, research focuses on developing these mappings, either focusing on specific hardware (e. g., neutral atom hardware [44]), or particular problems (knapsack problem [17], robot motion planning [15, 23]; see Table 2).
@@
Several quantum-inspired optimization strategies have emerged as byproducts of the research in quantum optimization. The marriage of generative modeling and quantum-inspired tensor network methods resulted in the generator-enhanced optimization (GEO) framework [48]. In GEO, a quantum-inspired generative model is used to propose new candidate solutions by identifying patterns in high-quality solutions. The performance of GEO has since been analyzed on the production planning use case, where it has been shown to perform on par with other conventionally used solvers [49, 50].
@@ Discussion
It is still unclear whether quantum computing can provide significant value for real-world optimization workloads [52]. With large, noisy quantum devices, the key challenge is identifying hardware-native problem instances, e. g., maximum independent set problems for neutral atom hardware that can be embedded with minimal overhead while remaining difficult for classical solvers [43, 53, 54]. While optimization of such problems on near-term quantum hardware devices has produced some impressive results [38], it is not yet feasible at the scales required for industrially relevant problems. However, with advancements in hardware and algorithms, scaling analyses suggest that quantum computing may eventually outperform classical methods for specific instances particularly suited to quantum devices [55, 56].
@@ 4 Simulation
HHL: The HHL algorithm uses quantum phase estimation (QPE) [83] to implicitly write the eigenvalues of matrix $A$ in a quantum register. It then stores the inverted eigenvalues in the quantum amplitudes, effectively inverting $A$. After reversing QPE and a postselection step, the algorithm prepares a state proportional to |$\ket{x}$, given that the input state was |$\ket{b}$. The overall complexity is $\mathcal{O}(\log(N)\kappa^{2}/\epsilon)$ [80], where $N$ corresponds to the dimension of $A$, $\kappa$ is its condition number quantifying the difficulty of inverting $A$, and $\epsilon$ is the precision of the solution. There is an exponential advantage in $N$ compared to the best-known classical methods. However, the promised speedup can be significantly reduced or even destroyed by considering the algorithmic overhead for state preparation, implementation of matrix oracles, or the readout of the full solution [84]. For example, in the case of FEM, the anticipated exponential speedup of the HHL algorithm is reduced to a polynomial speedup by considering the precision of the readout [85].
@@ 5 Materials Science and Quantum Chemistry
VQE: This type of algorithm starts with a variational quantum circuit (VQC) ansatz, which attempts to prepare the system’s ground state. In an iterative process, the variational parameters of the VQC are adapted via gradient descent, trying to minimize the system’s energy [13, 107]. VQE is considered a near-term algorithm that is suitable for NISQ computers. However, it suffers from trainability issues when large systems are considered or as the depth and complexity of the VQCs grow; see the barren plateau problem [77]. When a suitable set of parameters for the VQC is found and a trial state is prepared, the expected value of the Hamiltonian can be estimated. Thus, its energy and eigenstate can be found.
@@
Qubitization or Block encoding: All quantum circuits are unitary transformations. In this sense, unitary operators are native to quantum computers. There is no native way to encode arbitrary operators, e. g., general Hamiltonians or arbitrary operators, in a quantum register and operate over them. Such ability is necessary for quantum simulations. To simulate Hamiltonian dynamics using a quantum computer, it is necessary to use the technique of qubitization or block encoding [86], which is an FT algorithm primitive. The technique consists of encoding an arbitrary operator inside a larger unitary operator through auxiliary qubits. This technique is probabilistic, which adds to the sample complexity of the algorithms. When this is achieved, additional algorithms can perform computations with the encoded data, as we comment below.
@@ 6 Machine Learning
Progress in AI is driven by the scaling laws that describe the improvement of AI as a function of the progress in computing, data, and model parameters [119]. Similarly, Sutton’s ”bitter lesson” [120] emphasizes the long-term success of algorithms that leverage increased computational power over those relying on complex, domain-specific optimizations. As classical computing infrastructures hit scaling limits [121], quantum computing may eventually provide the computational capabilities needed to align with this principle.
Quantum machine learning (QML) [122, 123] integrates quantum algorithms with machine learning, aiming to reduce training times and enhance the scalability and quality of models. While some QML algorithms promise theoretical improvements in computational complexity [124, 125, 126, 127], the practical advantages of quantum machine learning have yet to be conclusively established on large-scale industry-relevant datasets. In the following sections, we describe near-term, fault-tolerant, and quantum-inspired QML, concluding with a discussion on the limitations of QML.
@@
Representing classical data on a quantum computer: QML algorithms can be applied to both quantum and classical data. For classical data, input data must be mapped to a quantum state using a procedure referred to as data or feature encoding. Notably, encoding methods that leverage entanglement, such as amplitude encoding, are particularly interesting because they allow for exponential data compression by encoding a feature vector of size $N$ into $\log N$ qubits. However, methods like amplitude encoding incur an exponential runtime in the number of qubits, and any potential quantum speed-up is lost at the encoding stage [84]. A promising solution lies in using approximations to address the challenge of exponential runtime in amplitude encoding. Jobst et al. [96] show that classical data with an algebraically decaying Fourier spectrum can be well-approximated by a quantum circuit with a linear number of nearest-neighbor two-qubit gates.
@@
Quantum kernel methods: Quantum kernel methods [128, 129] are one approach in QML for supervised learning tasks. They leverage quantum computing by replacing the classical kernel of a support vector machine (SVM) with a quantum kernel. In this approach, a classical data vector $\mathbf{x}$ is mapped into a high-dimensional feature space using a feature map that is represented by a parameterized quantum circuit $U(\mathbf{x})$. Quantum computers can efficiently evaluate the inner products between data points in this quantum feature space. By determining all pair-wise inner products of the embedded data points, the kernel matrix is generated. This kernel matrix is then used within the SVM framework to perform classification or regression tasks.
Supervised training of VQCs: Supervised training of VQCs [130, 131, 132] involves optimizing the parameters of a VQC to minimize a predefined cost function, typically derived from labeled data. The circuit processes input data $x$ encoded into quantum states by a feature map $U(x)$. This state is then evolved via a VQC $V(\theta)$, before the measurement is performed. The expectation value of the observable acts as the prediction of the model. The parameters $\theta$ are tuned during the training to minimize a cost function. A method for computing gradients on a quantum computer is the parameter-shift rule [130, 133].
Generative training of VQCs: Quantum states prepared by VQCs naturally represent probability distributions that can be used to generate synthetic data. Therefore, VQCs can also be used for generative learning. Instead of fitting the model’s output to the data label, the output is fitted to a distribution. Two widely used training routines in quantum generative modeling, as depicted in Figure 3, are the quantum circuit Born machine (QCBM) [134, 135] and the quantum generative adversarial network (QGAN) [136, 137], as depicted in Figure 4. For a comprehensive review of these methods, see the referenced literature [138]. The QCBM aims to minimize a statistical distance such as the Kullback-Leibler (KL) divergence, between the probability distributions of the generated and target data by adjusting model parameters with gradient-based or gradient-free optimization techniques.
@@ 7 Benchmarking
Aggregate benchmarks evaluate a larger part of the quantum computing stack, considering a broader range of attributes. An example is Quantum Volume (QV), which considers the number of qubits, the quality of gate operations, and the accuracy of measurements [171]. While the QV is a well-established benchmark, it has known drawbacks like its strong dependency on high-quality qubits, which favors technologies like ion-trapped quantum computers (which reach QV values of over a million [172, 173]) compared to superconducting hardware (currently reaching QV values well below 1000 [174]). Hence, other aggregate benchmarks like circuit layer operations per second (CLOPS) [175] or generalized volumetric benchmarks [176] based on QV address these shortcomings and offer alternatives.
@@ 8 Conclusion and Future Directions
Quantum Chemistry Fuel Cell [103] Quantinuum H1 using 6 qubits Simulation of Oxygen Reduction Reaction (ORR) on Platinum surface with active space of 2 electrons in 3 orbitals.
@@
Quantum hardware has significantly progressed. However, this progress has not matched initial industry expectations, particularly in scale. This limitation is evident from the data presented in Table 2, which summarizes our recent case studies. Our case studies utilized various quantum hardware platforms, e. g., approximately 5,000 qubits on D-Wave’s Advantage 4.1, 137 qubits on QuEra’s Aquilla, 11 qubits on IBM’s Kolkata, 6-10 qubits on Quantinuum’s H1 and IonQ’s Harmony machines.
@@
Future Directions: Hardware continues to advance, as evidenced by enhanced qubit fidelities [172], error correction breakthroughs [212, 213], and new distributed, multi-QPU architectures [214]. Emerging quantum-classical architectures, also referred to as HPCQC [215] or quantum-centric supercomputing [8], integrate Quantum Processing Units (QPUs) as accelerators into the high-performance computing stack. While these hardware trends are promising, they require adaptation to algorithms [216, 217, 218]. These adaptations often involve techniques such as circuit cutting [219], hardware-software co-design [207], and error mitigation to optimize performance and reliability.